The first equation is y=2x-8. If a second line passes through the points (0,4) and has a slope of -1, find the solution to the system.
Answer by mananth(16946)

The first equation is
If a second line passes through the points (0,4) and has a slope of -1, find the solution to the system.

Equation of line with slope m and passing through(x1,y1) is
(0,4) ~(x1,y1) and has a slope of m=-1
y-y1= m(x-x1)
y-4= -1(x-0)
y-4=-x
y=-x+4
substitute y
y=2x-8. (given)
-x+4=2x-8
3x=12
x=4
Therefore y = -x+4= 0
the solution to the system of equations is (4,0)
.